OpenStudy (anonymous):

38 Natashka134 hey could anybody help this hot young lady jk am not that hot jk again lol am so funny Which equation represents the relationship shown in the table below? x y 0 -3 1 -1 2 1 3 3 y = – x – 3 y = x – 3 y = 2x − 3 y = ­­–2x + 3

OpenStudy (anonymous):

for each 1 unit change in x, notice the y changes by 2...

OpenStudy (anonymous):

thank u hot owl

OpenStudy (anonymous):

aww... now i'm turning pink...

OpenStudy (anonymous):

can you take a stab at which one of those choices?

OpenStudy (p0sitr0n):

m = delta y / delta x = 3-(-3) / 3-0 = 6/3= 2 y = 2x+b guess its 3

OpenStudy (anonymous):

ya but u first

OpenStudy (anonymous):

-3+2.x

OpenStudy (anonymous):

thank u sweety